# Riemannian Optimization for Distance Geometry: A Study of Convergence, Robustness, and Incoherence

### v1

@@ Appendix B Restricted Isometry Results
where the first inequality is the triangle inequality, the second inequality is Cauchy-Schwarz, the third inequality is due to Lemma A.10 and (36), the fourth inequality is a result of (37) and (38), and the final inequality is due to (39). ∎
@@ Appendix C Local Convergence Results
We can now leverage the variational characterization of the spectral norm and local RIP, proven in Lemma B.12, to bound the following:
$-p^{2}(4\varepsilon_{0})\|\mathcal{P}_{\mathbb{T}}\bm{G}_{l}\|_{\mathrm{F}}^{2}\leq\left\langle\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l},\mathcal{P}_{\mathbb{T}_{l}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}-p^{2}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}\right\rangle\leq p^{2}(4\varepsilon_{0})\|\mathcal{P}_{\mathbb{T}}\bm{G}_{l}\|_{\mathrm{F}}^{2}.$
@@
Assume $\left\|\mathcal{P}_{\mathbb{T}_{l}}-p^{-2}\mathcal{P}_{\mathbb{T}_{l}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}\right\|\leq 4\varepsilon_{0}$ and $\alpha_{l}$ can be bounded as in Lemma C.1. Then the spectral norm of $\mathcal{P}_{\mathbb{T}_{l}}-\alpha_{l}\mathcal{P}_{\mathbb{T}_{l}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}$ can be bounded as
@@
where the first inequality comes from the triangle inequality, the second inequality comes from Local RIP in Lemma B.12, the third inequality comes from the stepsize bound in Lemma C.1, the fourth inequality again comes from Lemma B.12, and the remainder comes from algebraic simplification of terms. This finishes the proof. ∎
We can now prove Theorem 5.4.
@@
as $\bm{X}_{l+1}$ is the best $r$ approximation of $\bm{W}_{l}$. Plugging in $\bm{W}_{l}=\bm{X}_{l}+\alpha_{l}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}$, we see that
@@
It remains to bound each term individually. Using Lemma C.2, we see that
@@
Next, notice that from Lemma A.10 and the fact that $\mathcal{P}_{\mathbb{T}_{l}}\bm{X}_{l}=\bm{X}_{l}$,
@@
using Lemma A.10 and our initial local neighborhood assumption. Finally, we see that, following a similar argument as in the bound of $I_{2}$,
@@
where the second to last inequality follows from the same analysis conducted in Lemma B.12, just divided by 2. Collecting these results, we get
@@
where the first inequality follows from the triangle inequality and the second inequality follows from the fact that $\bm{W}_{0}$ is the best $r$ approximation of $\bm{X}_{0}$ by Eckart-Young-Mirsky[54]. We now need a bound for this last term. Notice that $\bm{W}_{0}-\bm{X}=\sum_{\bm{\alpha}\in\mathbb{I}}(p^{-1}\xi_{\bm{\alpha}}-1)\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle\bm{v}_{\bm{\alpha}}$ is a sum of zero-mean i.i.d random matrices, opening up use of Bernstein’s inequality. In order to use this, define $\bm{Z}_{\bm{\alpha}}=(p^{-1}\xi_{\bm{\alpha}}-1)\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle\bm{v}_{\bm{\alpha}}$. We need a bound on $\left\|\bm{Z}_{\bm{\alpha}}\right\|$ and $\left\|\mathbb{E}\left[\sum_{\bm{\alpha}\in\mathbb{I}}\bm{Z}_{\bm{\alpha}}\right]^{2}\right\|$. First, notice that
@@
where the second inequality comes from the fact that $p\leq 1$ and $\left\|\bm{v}_{\bm{\alpha}}\right\|<1$ from Lemma A.8. Next, notice that
@@
Now, as Lemma A.9, $\sum_{\bm{\alpha}}\bm{v}_{\bm{\alpha}}^{2}=\frac{n^{2}-2n+2}{4n}\bm{J}$. It follows that $\lambda_{\text{max}}\left(\sum_{\bm{\alpha}}\bm{v}_{\bm{\alpha}}^{2}\right)=\frac{n^{2}-2n+2}{4n}\leq\frac{n}{4}$ as $\bm{J}$ is an orthogonal projection matrix. Thus,
@@
verifying the probabilistic bound. To complete the proof, we use Lemma F.2, from which it follows that
@@ Appendix E Robustness Guarantees
In this section, we will prove Theorem 6.2. To begin, we will prove a result highlighting the dependencies of the size of the noise on the reconstruction of an object
@@
where the first inequality follows from the bound on $\|\bm{N}\|_{\infty}$, the second inequality follows from the definition of the Frobenius norm, and the final inequality follows from $\beta n^{1/2}r^{1/4}\log{n}>1$ and $\delta<1$. Now, notice that $\mathbb{E}\left[\bm{N}\bm{P}^{\top}\right]=\bm{0}$, and can be decomposed as the sum of independent random matrices as $\bm{N}\bm{P}^{\top}=\sum_{i=1}^{r}{\bm{n}^{i}}{\bm{p}^{i}}^{\top}$, where ${\bm{n}^{i}}$ and ${\bm{p}^{i}}$ are the $i$-th columns of $\bm{N}$ and $\bm{P}$, respectively. As such, we will use Theorem A.1. Now, using the bound on $\|\bm{N}\|_{\infty}$, we have that
@@
This result will follow from the classic Davis-Kahan $\sin\Theta$ Theorem, seen in Theorem A.4. Let $\mathbb{U},~\mathbb{\hat{U}}$ be the subspace spanned by the columns of $\bm{U},\hat{\bm{U}}$ respectively. First, as $\left\|\mathcal{P}_{U}-\mathcal{P}_{\hat{U}}\right\|_{\mathrm{F}}=\|\sin\Theta(\mathbb{U},\hat{\mathbb{U}})\|_{\mathrm{F}}$ from [95], we can see that
@@
where the third inequality follows from Lemma E.1 thus ending the proof. ∎
@@ Appendix F Incoherence Results
and as $\sum_{\bm{\alpha}\in\mathbb{I}}|H^{\bm{\alpha}\bm{\beta}}|\leq 2$ from Lemma A.8, the claim follows. An identical proof shows the second result, with $\mathcal{P}_{\mathbb{T}}$ in place of $\mathcal{P}_{U}$. ∎
@@
Let $\bm{X}\succeq\bm{0}$ be a $r$, $\nu$-incoherent matrix satisfying (20) with constant $\nu$. Then
$\left(\max_{\bm{\alpha}\in\mathbb{I}}\left|\langle\bm{X},\bm{w}_{\bm{\alpha}}\right\rangle|\right)\leq\frac{\nu r}{4n}\|\bm{X}\|$
@@
where the first inequality comes from the definition of the spectral norm, the penultimate line follows from a rescaled form of (13) in accordance with Assumption 5.1, and the final line follows from (20), thus concluding the proof. ∎
@@
Let $\mu$ be an a.s. bounded, mean-zero, sub-Gaussian distribution with positive definite covariance matrix $\bm{\Sigma}\in\mathbb{R}^{r\times r}$. Let $n$ points $\{\bm{p}_{i}\}_{i=1}^{n}\sim\mu$ be sampled i.i.d., and let $\bm{P}=[\bm{p}_{1}\dots\bm{p}_{n}]^{T}\in\mathbb{R}^{n\times d}$ be the corresponding point matrix with Gram matrix $\bm{X}$, which has condition number $\kappa$. Let $\|\bm{p}_{i}\|_{\psi_{2}}\leq K$ for some $K>0$. Then with probability at least $1-Cn^{-2}$ for some absolute constant $C>0$, the incoherence parameter of $\bm{X}$ is bounded by
@@
Next, following the argument of Lemma 3.3 but replacing $2r$ with $\mathbb{E}\left[\left(\bm{p}_{i}-\bm{p}_{j}\right)^{\top}\left(\bm{p}_{i}-\bm{p}_{j}\right)\right]$, we have that, with probability at least $1-Cn^{-2}$
@@
This gives us the bound $K\leq C\lambda_{1}(\bm{\Sigma})^{1/2}$ for some absolute constant $C>0$. Leveraging this, along with the fact that from Lemma 3.2 that $\lambda_{r}(\bm{X})\approx n\lambda_{r}(\bm{\Sigma})$, we have that for some $c>0$ with high probability that
@@
This concludes the proof.
@@ G.2 Riemannian Optimization
[…]
@@
Optimization over $\mathcal{N}_{r}$ has been investigated in detail for quite some time, and in particular retraction-based methods are of particular interest to this work [97, 98, 61, 99, 100, 51, 101, 102]. First-order retraction-based methodologies rely on the general principle of taking a descent step in the tangent space, followed by a retraction onto the manifold. In the case of first-order optimization on $\mathcal{N}_{r}$, the retraction map $\mathcal{H}_{r}$ is given by the hard thresholding operator, which is a thin spectral decomposition that takes $\bm{Y}=\sum_{i=1}^{n}\lambda_{i}\bm{u}_{i}\bm{u}_{i}^{\top}\mapsto\sum_{i=1}^{r}\lambda_{i}\bm{u}_{i}\bm{u}_{i}^{\top}$, where $|\lambda_{1}|\geq\cdots\geq|\lambda_{n}|$ are the ordered eigenvalues of $\bm{Y}$ and $\bm{u}_{i}$ are the corresponding eigenvectors of $\bm{Y}$.

### v2

@@ Appendix B Restricted Isometry Results
where the first inequality is the triangle inequality, the second inequality is Cauchy-Schwarz, the third inequality is due to A.10 and (37), the fourth inequality is a result of (38) and (39), and the final inequality is due to (40). ∎
@@ Appendix C Local Convergence Results
We can now leverage the variational characterization of the spectral norm and local RIP, proven in B.12, to bound the following:
$p^{2}(4\varepsilon_{0})\|\mathcal{P}_{\mathbb{T}}\bm{G}_{l}\|_{\mathrm{F}}^{2}\leq\left\langle\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l},\mathcal{P}_{\mathbb{T}_{l}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}-p^{2}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}\right\rangle\leq p^{2}(4\varepsilon_{0})\|\mathcal{P}_{\mathbb{T}}\bm{G}_{l}\|_{\mathrm{F}}^{2}.$
@@
Assume $\left\|\mathcal{P}_{\mathbb{T}_{l}}-p^{-2}\mathcal{P}_{\mathbb{T}_{l}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}\right\|\leq 4\varepsilon_{0}$ and $\alpha_{l}$ can be bounded as in C.1. Then the spectral norm of $\mathcal{P}_{\mathbb{T}_{l}}-\alpha_{l}\mathcal{P}_{\mathbb{T}_{l}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}$ can be bounded as
@@
where the first inequality comes from the triangle inequality, the second inequality comes from Local RIP in B.12, the third inequality comes from the stepsize bound in C.1, the fourth inequality again comes from B.12, and the remainder comes from algebraic simplification of terms. This finishes the proof. ∎
@@
as $\bm{X}_{l+1}$ is the best rank-$r$ approximation of $\bm{W}_{l}$. Plugging in $\bm{W}_{l}=\bm{X}_{l}+\alpha_{l}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}$, we see that
@@
It remains to bound each term individually. Using C.2, we see that
@@
Next, notice that from A.10 and the fact that $\mathcal{P}_{\mathbb{T}_{l}}\bm{X}_{l}=\bm{X}_{l}$,
@@
using A.10 and our initial local neighborhood assumption. Finally, we see that, following a similar argument as in the bound of $I_{2}$,
@@
where the second to last inequality follows from the same analysis conducted in B.12, just divided by 2. Collecting these results, we get
@@
where the first inequality follows from the triangle inequality and the second inequality follows from the fact that $\bm{W}_{0}$ is the best rank-$r$ approximation of $\bm{X}_{0}$ by Eckart-Young-Mirsky[57]. We now need a bound for this last term. Notice that $\bm{W}_{0}-\bm{X}=\sum_{\bm{\alpha}\in\mathbb{I}}(p^{-1}\xi_{\bm{\alpha}}-1)\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle\bm{v}_{\bm{\alpha}}$ is a sum of zero-mean i.i.d random matrices, opening up use of Bernstein’s inequality. In order to use this, define $\bm{Z}_{\bm{\alpha}}=(p^{-1}\xi_{\bm{\alpha}}-1)\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle\bm{v}_{\bm{\alpha}}$. We need a bound on $\left\|\bm{Z}_{\bm{\alpha}}\right\|$ and $\left\|\mathbb{E}\left[\sum_{\bm{\alpha}\in\mathbb{I}}\bm{Z}_{\bm{\alpha}}\right]^{2}\right\|$. First, notice that
@@
where the second inequality comes from the fact that $p\leq 1$ and $\left\|\bm{v}_{\bm{\alpha}}\right\|<1$ from A.8. Next, notice that
@@
Now, as A.9, $\sum_{\bm{\alpha}}\bm{v}_{\bm{\alpha}}^{2}=\frac{n^{2}-2n+2}{4n}\bm{J}$. It follows that $\lambda_{\text{max}}\left(\sum_{\bm{\alpha}}\bm{v}_{\bm{\alpha}}^{2}\right)=\frac{n^{2}-2n+2}{4n}\leq\frac{n}{4}$ as $\bm{J}$ is an orthogonal projection matrix. Thus,
@@
verifying the probabilistic bound. To complete the proof, we use F.2, from which it follows that
@@ Appendix E Robustness Guarantees
In this section, we will prove 6.2. To begin, we will prove a result highlighting the dependencies of the size of the noise on the reconstruction of an object
@@
where the first inequality follows from the bound on $\|\bm{N}\|_{\infty}$, the second inequality follows from the definition of the Frobenius norm, and the final inequality follows from $\beta n^{1/2}r^{1/4}\log{n}>1$ and $\delta<1$. Now, notice that $\mathbb{E}\left[\bm{N}\bm{P}^{\top}\right]=\bm{0}$, and can be decomposed as the sum of independent random matrices as $\bm{N}\bm{P}^{\top}=\sum_{i=1}^{r}{\bm{n}^{i}}{\bm{p}^{i}}^{\top}$, where $\bm{n}^{i}$} and $\bm{p}^{i}$} are the $i$-th columns of $\bm{N}$ and $\bm{P}$, respectively. As such, we will use A.1. Now, using the bound on $\|\bm{N}\|_{\infty}$, we have that
@@
This result will follow from the classic Davis-Kahan $\sin\Theta$ Theorem, seen in A.4. Let $\mathbb{U},~\mathbb{\hat{U}}$ be the subspace spanned by the columns of $\bm{U},\hat{\bm{U}}$ respectively. First, as $\left\|\mathcal{P}_{U}-\mathcal{P}_{\hat{U}}\right\|_{\mathrm{F}}=\|\sin\Theta(\mathbb{U},\hat{\mathbb{U}})\|_{\mathrm{F}}$ from [98], we can see that
@@
where the third inequality follows from E.1, thus ending the proof. ∎
@@ Appendix F Incoherence Results
and as $\sum_{\bm{\alpha}\in\mathbb{I}}|H^{\bm{\alpha}\bm{\beta}}|\leq 2$ from A.8, the claim follows. An identical proof shows the second result, with $\mathcal{P}_{\mathbb{T}}$ in place of $\mathcal{P}_{U}$. ∎
@@
Let $\bm{X}\succeq\bm{0}$ be a rank-$r$, $\nu$-incoherent matrix satisfying (21) with constant $\nu$. Then
$\left(\max_{\bm{\alpha}\in\mathbb{I}}\left|\langle\bm{X},\bm{w}_{\bm{\alpha}}\right\rangle|\right)\leq\frac{\nu r}{4n}\|\bm{X}\|.$
@@
where the first inequality comes from the definition of the spectral norm, the penultimate line follows from a rescaled form of (13) in accordance with 5.1, and the final line follows from (21), thus concluding the proof. ∎
@@
Let $\mu$ be an a.s. bounded, mean-zero, sub-Gaussian distribution with positive definite covariance matrix $\bm{\Sigma}\in\mathbb{R}^{r\times r}$. Let $n$ points $\{\bm{p}_{i}\}_{i=1}^{n}\sim\mu$ be sampled i.i.d., and let $\bm{P}=[\bm{p}_{1}\dots\bm{p}_{n}]^{\top}\in\mathbb{R}^{n\times d}$ be the corresponding point matrix with Gram matrix $\bm{X}$, which has condition number $\kappa$. Let $\|\bm{p}_{i}\|_{\psi_{2}}\leq K$ for some $K>0$. Then with probability at least $1-Cn^{-2}$ for some absolute constant $C>0$, the incoherence parameter of $\bm{X}$ is bounded by
@@
Next, following the argument of 3.3 but replacing $2r$ with $\mathbb{E}\left[\left(\bm{p}_{i}-\bm{p}_{j}\right)^{\top}\left(\bm{p}_{i}-\bm{p}_{j}\right)\right]$, we have that, with probability at least $1-Cn^{-2}$
@@
This gives us the bound $K\leq C\lambda_{1}(\bm{\Sigma})^{1/2}$ for some absolute constant $C>0$. Leveraging this, along with the fact that from 3.2 that $\lambda_{r}(\bm{X})\approx n\lambda_{r}(\bm{\Sigma})$, we have that for some $c>0$ with high probability that
@@
This concludes the proof. ∎
@@ G.2 Riemannian Optimization
The tangent space can be described as the set of all possible rank-up-to-$2r$ perturbations, represented as the sum of a perturbation in the column and row space, and is computed by looking at first-order perturbations of the spectral decomposition of $\bm{X}$[64]. Additionally, we can compute the orthogonal projection of an arbitrary $\bm{Y}\in\mathbb{R}^{n\times n}$ onto the tangent space at a point $T_{\bm{X}}\mathcal{N}_{r}$ as follows [64, 99, 54]:
@@
Optimization over $\mathcal{N}_{r}$ has been investigated in detail for quite some time, and retraction-based methods are of particular interest to this work [100, 101, 64, 102, 103, 54, 104, 105]. First-order retraction-based methodologies rely on the general principle of taking a descent step in the tangent space, followed by a retraction onto the manifold. In the case of first-order optimization on $\mathcal{N}_{r}$, the retraction map $\mathcal{H}_{r}$ is given by the hard thresholding operator, which is a thin spectral decomposition that takes $\bm{Y}=\sum_{i=1}^{n}\lambda_{i}\bm{u}_{i}\bm{u}_{i}^{\top}\mapsto\sum_{i=1}^{r}\lambda_{i}\bm{u}_{i}\bm{u}_{i}^{\top}$, where $|\lambda_{1}|\geq\cdots\geq|\lambda_{n}|$ are the ordered eigenvalues of $\bm{Y}$ and $\bm{u}_{i}$ are the corresponding eigenvectors of $\bm{Y}$.